OpenStudy (anonymous):

simplify completely x+2/7x^2-28

OpenStudy (mathstudent55):

You mean \( \dfrac{x + 2}{7x^2 - 28} \) ?

OpenStudy (anonymous):

yes idk how to writh them on the laptop with all the expressions lol

OpenStudy (mathstudent55):

The numerator is a binomial that is factored. The denominator is a binomial that can be factored. The first step is to factor the denominator completely.

OpenStudy (anonymous):

\[\frac{ 1 }{ 7x-14 } \]?

OpenStudy (anonymous):

is that it is that the answer

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

can someone help me please ANYBODY.

OpenStudy (mathstudent55):

I was trying to help you, but you didn't answer. Can you factor the denominator?

OpenStudy (anonymous):

NO I DID ANSWER IM SAYING THAT I DONT UNDERSTAND ALGEBRA IS ONE OF MY WEAK SUBJECTS

OpenStudy (anonymous):

I KNOW IT SHOULD B 4

OpenStudy (mathstudent55):

For the third time: Can you factor \(7x^2 - 28\) ?

OpenStudy (anonymous):

HOW DO U FACTOR WHAT R U SAYING I GUESS 4 WOULD BE THE FACTOR

OpenStudy (anonymous):

14

OpenStudy (mathstudent55):

The two terms 7x^2 and 28 have a common factor of 7, so you factor it out first: \( \dfrac{x + 2}{7(x^2 - 4)} \) Now you need to factor x^2 - 4 into two factors.

OpenStudy (mathstudent55):

\( =\dfrac{x + 2}{7(x + 2)(x - 2)}\) \( =\dfrac{\cancel{x + 2} ~1}{7(\cancel{x + 2})(x - 2)}\) \( =\dfrac{1}{7(x - 2)}\) \( =\dfrac{1}{7x - 14}\)